Plane A flies from airport A to airport B with 800 km/h. Plane B flies from airport B to airport A with 900 km/h. The distance between the 2 airports is 425 km. How long does it take before the 2 airplanes come across each other?
Answer by TimothyLamb(4379)

s = d/t
t = d/s
---
closure speed of the two airplanes = 800 + 900 = 1700 kph
t = 425/1700
---
answer:
t = 0.25 hours = 15 minutes
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php